OpenStudy (anonymous):

x quarts of milk cost g dollars, and y quarts of heavy cream cost g dollars. The price of one quarts of milk is what percent of the price of one quart of heavy cream? A)g B)y/x C)x/y D)100x/y E)100y/x

OpenStudy (anonymous):

anyone help?

OpenStudy (anonymous):

i would say either b or c...

OpenStudy (anonymous):

@Hope_nicole I thought so too, but the book says that the answer is e.. I just don't know how they got there haha

OpenStudy (anonymous):

ohh! i get it now!

OpenStudy (anonymous):

multiplying something by 100 is a percent...if its 93% that really means it is .93, so multiplying it by 100 makes it 93...does that make sense at all? lol

OpenStudy (anonymous):

unit price of milk is \(\frac{g}{x}\) and the unit price of cream is \(\frac{g}{y}\) ratio of unit price of milk to unit price of cream is \[\frac{\frac{g}{x}}{\frac{g}{y}}=\frac{y}{x}\]

OpenStudy (anonymous):

kk thanks tht helped a little)

OpenStudy (anonymous):

to turn that in to a percent, multiply by 100

OpenStudy (anonymous):

making your answer E

OpenStudy (anonymous):

sorry :/ i wish i could help more.

OpenStudy (anonymous):

@Hope_nicole @satellite73 thanks now i get it

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador